A galvanic cell Cr(s)|Cr2+(aq) || Cu2+(aq)|Cu(s) is constructed using a completely immersed Cr electrode that weighs 35.4 g and a Cu electrode immersed in 462 mL of 1.00 M Cu2+(aq) solution. A steady current of 0.0795 A is drawn from the cell as the electrons move from the Cr electrode to the Cu electrode.

(a) Which reactant is the limiting reactant in this cell?
Enter symbol


(b) How long does it take for the cell to be completely discharged?  s

(c) How much mass has the Cu electrode gained when the cell is completely discharged?  g

(d) What is the concentration of the Cu2+(aq) when the cell is completely discharged? (Assume that the limiting reactant is 100% reacted.)  M
Answer #1

A. Moles of Cu2+ = 462ml × 1 M = 0.462 moles

Moles of Cr = 35.4 /51.9961 g/mol = 0.680 moles

Cu2+ is limiting reagent

B. Moles of limiting reagent = Q/F = It / F

t = equivalent moles × F/ I = 2× 0.462 × 96485 / 0.0795

= 1121410.56 s

C. Mass of deposit = moles of Cu2+ × molar mass

= 0.462 × 63.546 = 29.35 g

D. Cu2+ = 0